OpenStudy (anonymous):

simplify-8^1/3/27^1/3

OpenStudy (anonymous):

\[\sqrt[3]{8}/\sqrt[3]{27}\]

OpenStudy (anonymous):

Since you have them both to the cube root you can write it like \[\sqrt[3]{8/27}\]